A long straight wire carries a current i. Let B1 be the magnetic field at a point P at a distance d from the wire. Consider a section of length l of this wire such that the point  P lies on a perpendicular bisector of the section. Let B2 be the magnetic field at this point due to this section only. Find the value of of d/l so that B2 differs from B1 by 1%.
 


Grade:10

1 Answers

ROSHAN MUJEEB
3 years ago

Solution :
B1=umoi2πdB1=umoi2πdand.B2=μoil4πd2×22–√B2=μoil4πd2×22
.Given.GivenB_2 - B_1 = B_1xx (1)/(100).⇒μ−−√oil4πd2=(1.01)μoi2πd⇒μoil4πd2=(1.01)μoi2πd
..(d)/(l) = (1.414)/(2.02) = 0.7` .